# Light-A-Video: Training-free Video Relighting via Progressive Light Fusion

### v1

@@ Abstract
Recent advancements in image relighting models, driven by large-scale datasets and pre-trained diffusion models, have enabled the imposition of consistent lighting. However, video relighting still lags, primarily due to the excessive training costs and the scarcity of diverse, high-quality video relighting datasets. A simple application of image relighting models on a frame-by-frame basis leads to several issues: lighting source inconsistency and relighted appearance inconsistency, resulting in flickers in the generated videos. In this work, we propose Light-A-Video, a training-free approach to achieve temporally smooth video relighting. Adapted from image relighting models, Light-A-Video introduces two key techniques to enhance lighting consistency. First, we design a Consistent Light Attention (CLA) module, which enhances cross-frame interactions within the self-attention layers to stabilize the generation of the background lighting source. Second, leveraging the physical principle of light transport independence, we apply linear blending between the source video’s appearance and the relighted appearance, using a Progressive Light Fusion (PLF) strategy to ensure smooth temporal transitions in illumination. Experiments show that Light-A-Video improves the temporal consistency of relighted video while maintaining the image quality, ensuring coherent lighting transitions across frames. Project page: https://bujiazi.github.io/light-a-video.github.io/.
Figure 1: Training-free video illumination control. Equipped with a pretrained image relighting model (e.g., IC-Light [63]) and a video diffusion model (e.g., CogVideoX [60] and AnimateDiff [13]), Light-A-Video enables training-free and zero-shot illumination control of any given video sequences or foreground sequences.
00footnotetext: * indicates equal contribution, † indicates corresponding author
@@ 1 Introduction
Illumination plays a crucial role in shaping our perception of visual content, impacting both its aesthetic quality and human interpretation of scenes. Relighting tasks [48, 33, 35, 66, 45, 18, 55, 67, 25], which focus on adjusting lighting conditions in 2D and 3D visual content, have long been a key area of research in computer graphics due to their broad practical applications, such as film production, gaming, and virtual environments. Traditional image relighting methods rely on physical illumination models, which struggle with accurately estimating real-world lighting and materials.
@@
To overcome these limitations, data-driven approaches [10, 40, 23, 65, 61, 20] have emerged, leveraging large-scale, diverse relighting datasets combined with pre-trained diffusion models. As the state-of-the-art image relighting model, IC-Light [63] modifies only the illumination of an image while maintaining its albedo unchanged. Based on the physical principle of light transport independence, IC-Light allows for controlled and stable illumination editing, such as adjusting lighting effects and simulating complex lighting scenarios. However, video relighting is significantly more challenging due to maintaining temporal consistency across frames. The scarcity of video lighting datasets and the cost of training further complicate the task. Thus, existing video relighting methods [65] struggle to deliver consistently high-quality results or are limited to specific domains, such as portraits.
Figure 2: Comparison of relighted frames using IC-Light and CLA module frame-by-frame. The line chart depicts the average optical flow intensity per frame. IC-Light shows a noticeable jitter between frames. Conversely, our CLA module achieves more stable light source generation over time, making the motion trend more consistent with the source video.
@@
In this work, we propose a training-free approach for video relighting, named Light-A-Video, which enables the generation of smooth, high-quality relighted videos without any training or optimization. As shown in Fig. 1, given a text prompt that provides a general description of the video and specified illumination conditions, our Light-A-Video pipeline can relight the input video in a zero-shot manner, fully leveraging the relighting capabilities of image-based models and the motion priors of the video diffusion model. To achieve this goal, we initially apply an image-relighting model to video-relighting tasks on a frame-by-frame basis, and observe that the generated lighting source is unstable across frames in the video. This instability leads to inconsistencies in the relighting of the objects’ appearances and significant flickering across frames. To stabilize the generated lighting source and ensure consistent results, we design a Consistent Light Attention (CLA) module within the self-attention layers of the image relighting model. As shown in Fig. 2, by incorporating additional temporally averaged features into the attention computation, CLA facilitates cross-frame interactions, producing a structurally stable lighting source. To further enhance the appearance stability across frames, we utilize the motion priors of the video diffusion model with a novel Progressive Light Fusion (PLF) strategy. Adhering to the physical principles of light transport, PLF progressively employs linear blending to integrate relighted appearances from the CLA into each original denoising target, which gradually guides the video denoising process toward the desired relighting direction. Finally, Light-A-Video serves as a complete end-to-end pipeline, effectively achieving smooth and consistent video relighting. As a training-free approach, Light-A-Video is not restricted to specific video diffusion models, making it highly compatible with a range of popular video generation backbones, including UNet-based and DiT-based models such as AnimateDiff [13], CogVideoX [60] and LTX-Video [15]. Our contributions are summarized as follows:
  • We present Light-A-Video, the first training-free video relighting model, which leverages the capabilities of image relighting models to generate high-quality, temporally consistent relighted videos.
@@
  • An end-to-end pipeline is proposed with two key designs: the Consistent Light Attention, which enhances the stability of lighting source across frames, and the Progressive Light Fusion, which gradually injects lighting information to ensure temporal consistency for video appearance.
  • Experiments show the effectiveness of our method across multiple settings. It not only supports relighting of the entire input video, but also supports relighting the input foreground sequences with background generation.
@@ 2 Related Work
Video diffusion models [3, 5, 6, 13, 51, 54, 17, 60, 2, 62, 4] aim to synthesize temporally consistent image frames based on provided conditions, such as a text prompt or an image prompt. In the realm of text-to-video (T2V) generation, the majority of methods [51, 13, 5, 6, 3, 62] train additional motion modeling modules from existing text-to-image architectures to model the correlation between video frames, while others [17, 60] train from scratch to learn video priors. For image-to-video (I2V) tasks that enhance still images with reasonable motions, a line of research [57, 7] proposes novel frameworks dedicated to image animation. Some approaches [12, 64, 14] serve as plug-to-play adapters. Stable Video Diffusion [2] fine-tune pre-trained T2V models for I2V generation, achieving impressive performance. Numerous works [34, 30, 27] focus on controllable generation, providing more controllability for users. Video diffusion models, due to their inherent video priors, are capable of synthesizing smooth and consistent video frames that are both content-rich and temporally harmonious.
2.2 Learning-based Illumination Control.
@@
Over the past few years, a variety of lighting control techniques [48, 33, 35] for 2D and 3D visual content based on deep neural networks have been proposed, especially in the field of portrait lighting [47, 1, 44, 46, 23], along with a range of baselines [66, 45, 18, 55, 67] aimed at improving the effectiveness, accuracy, and theoretical foundation of illumination modeling. Recently, owing to the rapid development of diffusion-based generative models, a number of lighting control methods [40, 61, 10, 20] utilizing diffusion models have also been introduced. Relightful Harmonization [40] focuses on harmonizing sophisticated lighting effects for the foreground portrait conditioning on a given background image. SwitchLight [23] suggests training a physically co-designed framework for human portrait relighting. IC-Light [63] is a state-of-the-art approach for image relighting. LumiSculpt [65] enables consistent lighting control in T2V generation models for the first time. However, in the domain of video lighting, the aforementioned approaches fail to simultaneously ensure precise lighting control and exceptional visual quality. This work incorporates a pre-trained image lighting control model into the denoising process of a T2V model through progressive guidance, leveraging the latter’s video priors to facilitate the smooth transfer of image lighting control knowledge, thereby enabling accurate and harmonized control of video lighting.
2.3 Video Editing with Diffusion Models.
@@
In recent years, diffusion-based video editing has undergone significant advancements. Some researches [29, 53, 56, 30, 32] adopt pretrained text-to-image (T2I) backbones for video editing. Another line of approaches [59, 58, 8, 19] leverages pre-trained optical flow models to enhance the temporal consistency of output video. Numerous studies [38, 11, 21] have concentrated on exploring zero-shot video editing approaches. COVE [52] leverages the inherent diffusion feature correspondence proposed by DIFT [49] to achieve consistent video editing. SDEdit [31] utilizes the intrinsic capability of diffusion models to refine details based on a given layout, enabling efficient editing for both image and video. Despite the remarkable performance of existing video editing techniques in various settings, there remains a lack of approaches specifically designed for controlling the lighting of videos.
@@ 3.1 Diffusion Model
$\mathbf{v}=\sqrt{\bar{\alpha}_{t}}\epsilon-\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}_{0}$ (3)
@@
$\hat{\mathbf{z}}_{0\leftarrow t}$ represents the denoising target based on $\mathbf{v}_{t}$.
@@ 4 Light-A-Video
In Section 4.1, we define the objectives of the video relighting task. In Section 4.2, we reveal that per-frame image relighting for video sequence suffers from lighting source inconsistency and accordingly propose Consistent Lighting Attention (CLA) module for enhanced lighting stability in per-frame image relighting results. In Section 4.3, we provide a detailed description of the proposed Light-A-Video pipeline, which achieves stable lighting source generation and temporally consistent relighted video through Progressive Light Fusion (PLF) strategy.
@@ 4.1 Problem Formulation
Given a source video and a lighting condition $c$, the objective of video relighting is to render the source video into the relighted video that aligns with the given condition $c$. Unlike image relighting solely concentrates on appearance, video relighting raises extra challenges in maintaining temporal consistency and motion preservation, necessitating high-quality visual coherence across frames.
@@ 4.2 Consistent Light Attention
To improve inter-frame information integration and generate a stable light source, we propose a Consistent Light Attention (CLA) module. Specifically, for each self-attention layer in the IC-Light model, a video feature map $\mathbf{h}\in\mathbb{R}^{(b\times f)\times(h\times w)\times c}$ serves as the input, where $b$ is the batch size and $f$ is the number of video frames, $h$ and $w$ denote the height and width of the feature map, with $h\times w$ representing the number of tokens for attention computation. With linearly projections, $\mathbf{h}$ is projected into query, key and value features $Q,K,V\in\mathbb{R}^{(b\times f)\times(h\times w)\times c}$. The attention computation is defined as follows:
$\operatorname{Self-Attn}(Q,K,V)=\operatorname{Softmax}\left(\frac{QK^{T}}{\sqrt{c}}\right)V$
Note that the naive method treats the frame dimension as the batch size, performing self-attention frame by frame with the image relighting model, which results in each frame attending only to its own features. For our CLA module as shown in Fig. 3, a dual-stream attention fusion strategy is applied. Given the input $\mathbf{h}$, the original stream directly feeds the feature map into the attention module to compute frame-by-frame attention, resulting in the output $\mathbf{h}^{\prime}_{1}$. The average stream first reshapes $\mathbf{h}$ into $\mathbb{R}^{b\times f\times(h\times w)\times c}$, averages it along the temporal dimension, then expands it $f$ times to obtain $\mathbf{\bar{h}}$. Specifically, the average stream mitigates high-frequency temporal fluctuations, thereby facilitating the generation of a stable background light source across frames. Meanwhile, the original stream retains the original high-frequency details, thereby compensating for the detail loss incurred by the averaging process. Then, $\mathbf{\bar{h}}$ is input into the self-attention module and the output is $\mathbf{\bar{h}}^{\prime}_{2}$. The final output $\mathbf{h}_{o}^{\prime}$ of the CLA module is a weighted average between two streams, with the trade-off parameter $\gamma$,
$\mathbf{h}_{o}^{\prime}=(1-\gamma)\mathbf{h}^{\prime}_{1}+\gamma\mathbf{\bar{h}}^{\prime}_{2}$ (7)
With the help of CLA, the result can capture global context across the entire video, and generated a more stable relighting result as shown in Fig. 2.
@@ 4.3 Progressive Light Fusion
CLA module improves cross-frame consistency but lacks pixel-level constraints, leading to inconsistencies in appearance details. To address this, we leverage motion priors in Video Diffusion Model (VDM), which are trained on large-scale video datasets and use temporal attention module to ensure consistent motion and lighting changes. The novelty of our Light-A-Video lies in progressively injecting the relighting results as guidance into the denoising process.
Figure 4: Visualization of the PLF Strategy. During the denoising process of the VDM, the PLF strategy progressively replaces the original Consistent Target $P^{t}_{c}$ with the Fusion Target $P_{f}^{t}$, guiding the denoising direction from $\mathbf{z}^{v}_{0\leftarrow t}$ to $\tilde{\mathbf{v}}_{t}$.
@@
Driven by the motion priors in the VDM, the denoising process encourage $\hat{\mathbf{z}}_{0\leftarrow t}$ to be temporally consistent. Thus, we define this target as the video Consistent Target $\mathbf{z}^{v}_{0\leftarrow t}$ with environment illumination $L^{v}_{t}$. However, discrepancies still exist between the predicted $\mathbf{z}^{v}_{0\leftarrow t}$ and the original video. To address this issue, we incorporate the differences as details compensation into the Consistent Target at each step, thereby enhancing detail consistency between the relighted video and the original video. Then, $\mathbf{z}^{v}_{0\leftarrow t}$ is sent into the CLA module to obtain the relighted latent, which serves as the Relight Target $\mathbf{z}^{r}_{0\leftarrow t}$ with the illumination $L^{r}_{t}$ for the $t$-th denoising step. Aligning with the light transport theory in Section 3.2, a pre-trained VAE $\{{\mathcal{E}(\cdot)},{\mathcal{D}(\cdot)}\}$ is used to decode the two targets into pixel level, yielding the image appearances $\mathbf{I}^{v}_{t}=\mathcal{D}(\mathbf{z}^{v}_{0\leftarrow t})$ and $\mathbf{I}^{r}_{t}=\mathcal{D}(\mathbf{z}^{r}_{0\leftarrow t})$, respectively. Refer to Eq. 6, the fusing appearance $\mathbf{I}^{f}_{t}$ can be formulated as,
@@
We observe that directly using encoded latent $\mathcal{E}(\mathbf{I}^{f}_{t})$ as the new target at each step results in suboptimal performance. The reason is that the gap between the two targets is excessively large, surpassing the refinement capability of the VDM and causing visible temporal lighting jitter. To mitigate this gap, a progressive lighting fusion strategy is proposed. Specifically, a fusion weight $\lambda_{t}$ is introduced, which decreases as denoising progresses, gradually reducing the influence of the relight target. The progressive light fusion appearance is defined as $\mathbf{I}^{p}_{t}$,
@@
The encoded latent $\tilde{\mathbf{z}}_{0\leftarrow t}=\mathcal{E}(\mathbf{I}^{p}_{t})$ is utilized as the Fusion Target for step $t$, replacing the original $\hat{\mathbf{z}}_{0\leftarrow t}$. Based on the fusion target, the less noisy latent $\mathbf{z}_{t-1}$ can be computed with DDIM scheduler with $v$-prediction:
$a_{t}=\sqrt{\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}},\\ b_{t}=\sqrt{\bar{\alpha}_{t-1}}-\sqrt{\bar{\alpha}_{t}}a_{t}$ (10)
$\mathbf{z}_{t-1}=a_{t}\mathbf{z}_{t}+b_{t}\tilde{\mathbf{z}}_{0\leftarrow t}$ (11)
@@
which means our PLF strategy essentially refines $\mathbf{v}_{t}$ iteratively and guides the denoising process towards the relighting direction shown in Fig 4. Notably, other schedulers (e.g., Euler Scheduler [22], Flow Matching [28], etc.) with predicted noise-free components for each step are also applicable. Finally, as the denoising progresses, we achieve a smooth and consistent injection of illumination, ensuring coherent video relighting.

### v2

@@ Abstract
Recent advancements in image relighting models, driven by large-scale datasets and pre-trained diffusion models, have enabled the imposition of consistent lighting. However, video relighting still lags, primarily due to the excessive training costs and the scarcity of diverse, high-quality video relighting datasets. A simple application of image relighting models on a frame-by-frame basis leads to several issues: lighting source inconsistency and relighted appearance inconsistency, resulting in flickers in the generated videos. In this work, we propose Light-A-Video, a training-free approach to achieve temporally smooth video relighting. Adapted from image relighting models, Light-A-Video introduces two key techniques to enhance lighting consistency. First, we design a Consistent Light Attention (CLA) module, which enhances cross-frame interactions within the self-attention layers of the image relight model to stabilize the generation of the background lighting source. Second, leveraging the physical principle of light transport independence, we apply linear blending between the source video’s appearance and the relighted appearance, using a Progressive Light Fusion (PLF) strategy to ensure smooth temporal transitions in illumination. Experiments show that Light-A-Video improves the temporal consistency of relighted video while maintaining the relighted image quality, ensuring coherent lighting transitions across frames. Project page: https://bujiazi.github.io/light-a-video.github.io/.
Figure 1: Training-free video relighting. Equipped with an image relighting model (e.g., IC-Light [62]) and a video diffusion model (e.g., CogVideoX [59] and AnimateDiff [13]), Light-A-Video enables training-free video relighting for given video sequences or foreground sequences.
@@ 1 Introduction
Illumination plays a crucial role in shaping our perception of visual content, impacting both its aesthetic quality and human interpretation of scenes. Relighting tasks [47, 32, 34, 65, 44, 17, 54, 66, 24], which focus on adjusting lighting conditions in 2D and 3D visual content, have long been a key area of research in computer graphics due to their broad practical applications, such as film production, gaming, and virtual environments. Traditional image relighting methods, which rely on physical illumination models, are typically confined to controlled laboratory settings and struggle to generalize to complex, unconstrained real-world lighting and material estimation.
@@
In order to address these limitations, data-driven approaches [10, 39, 22, 64, 60, 19] have emerged, leveraging large-scale, diverse relighting datasets combined with pre-trained diffusion models. As the state-of-the-art image relighting model, IC-Light [62] modifies only the illumination of an image while maintaining its albedo unchanged. Based on the physical principle of light transport independence, IC-Light allows for controllable and stable illumination editing, such as adjusting lighting effects and simulating complex lighting scenarios. However, video relighting is significantly more challenging as it needs to maintain temporal consistency across frames. The scarcity of video lighting datasets and the high training costs further complicate the task. Thus, existing video relighting methods [64] struggle to deliver consistently high-quality results or are limited to specific domains, such as portraits.
Figure 2: Relighted frames of vanilla IC-Light and “IC-Light + CLA” . The line chart depicts the average optical flow intensity between adjacent frames. Since IC-Light performs image relighting based on each independent frame, its results show a noticeable jitter between frames, especially in the generated background lighting. Conversely, the proposed CLA facilitates consistent lighting generation by forcing interaction between frames.
@@
In this work, we propose a training-free approach for video relighting, named Light-A-Video, which enables the generation of smooth, high-quality relighted videos without any additional training or optimization. As shown in Fig. 1, given a text prompt that provides a general description of the video and specified illumination conditions, our Light-A-Video pipeline can relight the input video in a zero-shot manner, fully leveraging the relighting capabilities of image-based models and the motion priors of the video diffusion model. To achieve this goal, we initially apply an image-relighting model to video-relighting tasks on a frame-by-frame basis, and observe that the generated lighting source is unstable across the video frames. This instability leads to inconsistencies in the relighting of the objects’ appearances and significant flickering across frames. To stabilize the generated lighting source and ensure consistent results, we design a Consistent Light Attention (CLA) module within the self-attention layers of the image relighting model. As shown in Fig. 2, by incorporating additional temporally averaged features into the attention computation, CLA facilitates cross-frame interactions, producing a structurally stable lighting source. To further enhance the appearance stability across frames, we utilize the motion priors of the video diffusion model with a novel Progressive Light Fusion (PLF) strategy. Adhering to the physical principles of light transport, PLF progressively employs linear blending to integrate relighted appearances from the CLA into each original denoising target, which gradually guides the video denoising process toward the desired relighting direction. Finally, Light-A-Video serves as a complete end-to-end pipeline, effectively achieving smooth and consistent video relighting. As a training-free approach, Light-A-Video is not restricted to specific video diffusion models, making it highly compatible with a range of popular video generation backbones, including UNet-based and DiT-based models such as AnimateDiff [13] and CogVideoX [59]. Our contributions are summarized as follows:
  • We propose Light-A-Video, a novel training-free video relighting framework that generalizes the capabilities of image relighting models to the video domain, enabling flexible and temporally consistent video relighting.
@@
  • We introduce two key designs: a consistent light attention module to enhance the stability of lighting sources across frames, and a progressive light fusion strategy gradually injects lighting information to facilitate temporal consistency in video appearance.
  • Extensive experiments under various scenarios demonstrate the effectiveness and versatility of the proposed method, which not only supports relighting the entire video sequences but also enables relighting for given foreground sequences.
@@ 2 Related Work
Video Diffusion Models. Video diffusion models [3, 5, 6, 13, 50, 53, 16, 59, 2, 61, 4] aim to synthesize temporally consistent image frames based on provided conditions, such as a text prompt or an image prompt. In the realm of text-to-video (T2V) generation, the majority of methods [50, 13, 5, 6, 3, 61] train additional motion modeling modules from existing text-to-image architectures to model the correlation between video frames, while others [16, 59] train from scratch to learn video priors. For image-to-video (I2V) tasks that enhance still images with reasonable motions, a line of research [56, 7] proposes novel frameworks dedicated to image animation. Some approaches [12, 63, 14] serve as plug-to-play adapters. Stable Video Diffusion [2] fine-tune pre-trained T2V models for I2V generation, achieving impressive performance. Numerous works [33, 29, 26] focus on controllable generation, providing more controllability for users. Video diffusion models, due to their inherent video priors, are capable of synthesizing smooth and consistent video frames that are both content-rich and temporally harmonious.
@@
Learning-based Illumination Control. Over the past few years, a variety of lighting control techniques [47, 32, 34] for 2D and 3D visual content based on deep neural networks have been proposed, especially in the field of portrait lighting [46, 1, 43, 45, 22], along with a range of baselines [65, 44, 17, 54, 66] aimed at improving the effectiveness, accuracy, and theoretical foundation of illumination modeling. Recently, owing to the rapid development of diffusion-based generative models, a number of lighting control methods [39, 60, 10, 19] utilizing diffusion models have also been introduced. Relightful Harmonization [39] focuses on harmonizing sophisticated lighting effects for the foreground portrait conditioning on a given background image. SwitchLight [22] suggests training a physically co-designed framework for human portrait relighting. IC-Light [62] is a state-of-the-art approach for image relighting. LumiSculpt [64] enables consistent lighting control in T2V generation models for the first time. However, in the domain of video lighting, the aforementioned approaches fail to simultaneously ensure precise lighting control and exceptional visual quality. This work incorporates a pre-trained image lighting control model into the denoising process of a T2V model through progressive guidance, leveraging the latter’s video priors to facilitate the smooth transfer of image lighting control knowledge, thereby enabling accurate and harmonized control of video lighting.
@@
Video Editing with Diffusion Models. In recent years, diffusion-based video editing has undergone significant advancements. Some researches [28, 52, 55, 29, 31] adopt pre-trained text-to-image (T2I) backbones for video editing. Another line of approaches [58, 57, 8, 18] leverages pre-trained optical flow models to enhance the temporal consistency of output video. Numerous studies [37, 11, 20] have concentrated on exploring zero-shot video editing approaches. COVE [51] leverages the inherent diffusion feature correspondence proposed by DIFT [48] to achieve consistent video editing. SDEdit [30] utilizes the intrinsic capability of diffusion models to refine details based on a given layout, enabling efficient editing for both image and video. Despite the remarkable performance of existing video editing techniques in various settings, there remains a lack of approaches specifically designed for controlling the lighting of videos.
@@ 3.1 Diffusion Model
$\mathbf{v}=\sqrt{\bar{\alpha}_{t}}\epsilon-\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}_{0}.$ (3)
@@
$\hat{\mathbf{z}}_{0\leftarrow t}$ represents the denoising target at time step $t$.
@@ 4 Light-A-Video
Section 4.1 defines the objectives of the video relighting task. Section 4.2 reveals that per-frame image relighting for video sequences suffers from lighting source inconsistency and accordingly proposes the Consistent Lighting Attention (CLA) module for enhanced stability in generated lighting source across frames. Section 4.3 represents the Progressive Light Fusion (PLF) strategy for temporally consistent video appearance generation.
@@ 4.1 Problem Formulation
Given a source video and a lighting condition $c$, the objective of video relighting is to render the source video into the relighted video that maintains the motion in the source video while aligning the lighting condition $c$. Unlike image relighting that solely concentrates on appearance, video relighting raises extra challenges in maintaining temporal consistency and motion preservation, necessitating high-quality visual coherence across frames.
@@ 4.2 Consistent Light Attention
To improve inter-frame information integration and generate a stable light source, we propose a Consistent Light Attention (CLA) module. Specifically, for each self-attention layer in the IC-Light model, a video feature map $\mathbf{h}\in\mathbb{R}^{(b\times f)\times(h\times w)\times d}$ serves as the input, where $b$ is the batch size and $f$ is the number of video frames, $h$ and $w$ denote the height and width of the feature map, with $h\times w$ representing the number of tokens for attention computation. With linearly projections, $\mathbf{h}$ is projected into query, key and value features $Q,K,V\in\mathbb{R}^{(b\times f)\times(h\times w)\times d}$. The attention computation is defined as follows:
$\operatorname{Self-Attn}(Q,K,V)=\operatorname{Softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V.$ (7)
Note that the naive method treats the frame dimension as the batch size, performing self-attention frame by frame with the image relighting model, which results in each frame attending only to its features. For the CLA module, as shown in Fig. 3, a dual-stream attention fusion strategy is applied. Given the input feature $\mathbf{h}$, the original stream directly feeds the feature map into the attention module to compute frame-by-frame attention, resulting in the output $\mathbf{h}^{\prime}_{1}$. The average stream first reshapes $\mathbf{h}$ into $\mathbb{R}^{b\times f\times(h\times w)\times d}$, averages it along the temporal dimension, then expands it $f$ times to obtain $\mathbf{\bar{h}}$. Specifically, the average stream mitigates high-frequency temporal fluctuations, thereby facilitating the generation of a stable background light source across frames. Meanwhile, the original stream retains the original high-frequency details, thereby compensating for the detail loss incurred by the averaging process. Then, $\mathbf{\bar{h}}$ is input into the self-attention module and the output is $\mathbf{\bar{h}}^{\prime}_{2}$. The final output $\mathbf{h}_{o}^{\prime}$ of the CLA module is a weighted average between two streams, with the trade-off parameter $\gamma$,
$\mathbf{h}_{o}^{\prime}=(1-\gamma)\mathbf{h}^{\prime}_{1}+\gamma\mathbf{\bar{h}}^{\prime}_{2}.$ (8)
With the help of CLA, the result can capture global context across the entire video and generate a more stable lighting source, as shown in Fig. 2.
@@ 4.3 Progressive Light Fusion
CLA module improves cross-frame consistency but lacks pixel-level constraints, leading to inconsistencies in appearance details. To address this, we leverage motion priors in the Video Diffusion Model (VDM), which are trained on large-scale video datasets and use a temporal attention module to ensure consistent motion and lighting changes. The novelty of our Light-A-Video lies in progressively injecting the relighting results as guidance into the denoising process.
@@
Figure 4: Visualization of the PLF strategy. During the denoising process of the VDM, the PLF strategy progressively replaces the original Consistent Target $\mathbf{z}^{v}_{0\leftarrow t}$ with the Fusion Target $\tilde{\mathbf{z}}_{0\leftarrow t}$, guiding the denoising direction from $\mathbf{v}_{t}$ to $\tilde{\mathbf{v}}_{t}$.
Figure 5: Visualization of the detail compensation. $\Delta d_{m}$ records the difference between $\hat{\mathbf{z}}_{0\leftarrow m}$ and the source video in the first denoising step, which is used as a detail compensation component for detail preservation in the consistent target.
Driven by the motion priors in the VDM, the denoising process encourage $\hat{\mathbf{z}}_{0\leftarrow t}$ to be temporally consistent. Thus, we define this target as the video Consistent Target $\mathbf{z}^{v}_{0\leftarrow t}$ with environment illumination $L^{v}_{t}$. However, discrepancies still exist between the predicted $\hat{\mathbf{z}}_{0\leftarrow t}$ and the original video, resulting in detail loss in the relighted video. To address this issue, as shown in Fig. 5, details compensation $\Delta d_{m}$ is incorporated into the $\mathbf{z}^{v}_{0\leftarrow t}$ at each step. Then, $\mathbf{z}^{v}_{0\leftarrow t}$ is sent into the CLA module to obtain the relighted latent, which serves as the Relight Target $\mathbf{z}^{r}_{0\leftarrow t}$ with the illumination $L^{r}_{t}$ for the $t$-th denoising step. Aligning with the light transport theory in Section 3.2, a pre-trained VAE $\{{\mathcal{E}(\cdot)},{\mathcal{D}(\cdot)}\}$ is used to decode the two targets into pixel level, yielding the image appearances $\mathbf{I}^{v}_{t}=\mathcal{D}(\mathbf{z}^{v}_{0\leftarrow t})$ and $\mathbf{I}^{r}_{t}=\mathcal{D}(\mathbf{z}^{r}_{0\leftarrow t})$, respectively. Refer to Eq. 6, the fusing appearance $\mathbf{I}^{f}_{t}$ can be formulated as:
@@
It is observed that directly using encoded latent $\mathcal{E}(\mathbf{I}^{f}_{t})$ as the new target at each step results in suboptimal performance. This is attributed to the excessively large gap between the two targets, which exceeds the refinement capability of the VDM and consequently causes visible temporal lighting jitter. To mitigate this gap, a progressive lighting fusion strategy is proposed. Specifically, a fusion weight $\lambda_{t}$ is introduced, which decreases as denoising progresses, thereby gradually reducing the influence of the relight target. The progressive light fusion appearance is defined as $\mathbf{I}^{p}_{t}$, i.e.,
@@
The encoded latent $\tilde{\mathbf{z}}_{0\leftarrow t}=\mathcal{E}(\mathbf{I}^{p}_{t})$ is utilized as the Fusion Target for step $t$, replacing the original $\mathbf{z}^{v}_{0\leftarrow t}$. Based on the fusion target, the less noisy latent $\mathbf{z}_{t-1}$ can be computed with DDIM scheduler with $v$-prediction:
$a_{t}=\sqrt{\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}},\\ b_{t}=\sqrt{\bar{\alpha}_{t-1}}-\sqrt{\bar{\alpha}_{t}}a_{t},$ (11)
$\mathbf{z}_{t-1}=a_{t}\mathbf{z}_{t}+b_{t}\tilde{\mathbf{z}}_{0\leftarrow t}.$ (12)
@@
which means PLF essentially refines $\mathbf{v}_{t}$ iteratively and guides the denoising process towards the relighting direction, as shown in Fig 4. Other schedulers capable of modeling the denoising direction, such as Euler Scheduler [21] and Rectified Flow [27], are also applicable. As the denoising progresses, smooth and consistent illumination injection is achieved, ensuring coherent video relighting.